OpenStudy (anonymous):

HELP PLEASE I NEED HELP IS THIS ANSWER CORRECT OR NOT???????????????? [ 62−( 24⋅30 ) ]⋅[4− (−3) 2 ] i got -100 is it correct????

OpenStudy (anonymous):

ok reverifying..

OpenStudy (anonymous):

I think not

OpenStudy (anonymous):

My calculator tells me it's -6580

OpenStudy (anonymous):

ok why?

OpenStudy (anonymous):

essentially this is -658*10

OpenStudy (anonymous):

ok how tho?

OpenStudy (anonymous):

Ok. let's dissect the equation

OpenStudy (anonymous):

very correct @Robert136

OpenStudy (anonymous):

[62-(24*30)]=62-720=-658 right?

OpenStudy (anonymous):

Now look at left [4-(-3)2]=10

OpenStudy (anonymous):

Order of operation is the first thing you need to bear in mind

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I hope this helps...

OpenStudy (anonymous):

You can throw me any doubts

OpenStudy (anonymous):

Ok your inaction seems to imply disposal of this issue. Good luck;)